OpenStudy (anonymous):

If V varies as the cube of T and V=16, when T is 4 find the constant of the variation

OpenStudy (anonymous):

V = kT^3 16 = k*4^3 16 = k * 64 1/4 = k

OpenStudy (anonymous):

If A varies jointly as the square of B and the square of C and inversely as the cube of D and A=18 when b=4 c=3 d=2